Tia, Zarko, Backfish, Carrie, Keith, Corey, Xiao, Sam, Tom
- Ramp a couple times today.
- Earliest protons to BNB is Wed/Thur.
- Studies during day hours, running at night.
- Permit sign off is tomorrow.
- Checking water resistivity needed to run power supplies.
- Measured 7MΩ this morning, may need more resistivity, will find out.
- Fixing a controlled access gate.
- Half shells came in, need to drill and fit them.
- Ask MCR what the run plan is so we can run untriggered data for timing measurements.
- Computer network in LArTF is refreshed.
- Josh wants to run for a short while on neutrinos ~1 month, then remove it from hall.
- Waiting to replace a water pump.
- Doing DAQ upgrades & testing on cosmics while waiting.